• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jin Huang

    Log entry time 03:32:32 on July 01, 2010

    Entry number 329242

    This entry is a followup to: 328374

    keyword=Optics Database Update

    Optics database has been updated to show clear foil reconstruction.

    More carbon foils has been missing since PREX running, which made identifying foils difficult. By studying central ray of PREX data (taken in March, when there were still 5 carbon foils), z position of Ta foil was measured as 4.4+-0.5(stat)mm upstream to central optics foil. Together with current data, this strongly suggests that central optics foil has fallen out, which is consitent with Seamus's study. (Halog 329207). It will be more clear after openning target chamber today.

    With Vince's calibration of beam position, a quick calibration on VZ has been performed, giving an online resolution of 2cm on z or 2mm on tg_y.

    Online replay database has been updated accordingly.

    A copy of this log entry has been emailed to: riordan, vasulk, saenboon, yqiang

    vertex z for run 1212+1214 (Combinaiton of sieve out Ta and Optics target run, dp = 0%, 4 peaks are 3 carbon + 1 Ta foil)